OpenStudy (xoterrixo13):

Find the distance between the two points. Round to the nearest tenth if necessary. (2, 5) and (–1, –5)

hartnn (hartnn):

Distance between points (x1,y1) and (x2,y2) is \(\huge d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\)

OpenStudy (xoterrixo13):

1 109 10.4 3

hartnn (hartnn):

just plug in points.

OpenStudy (xoterrixo13):

3

OpenStudy (anonymous):

lol..Hartann u r Fast....)

OpenStudy (xoterrixo13):

3

hartnn (hartnn):

nopes, \(\sqrt{10^2+3^2}=...?\)

OpenStudy (xoterrixo13):

10.4

hartnn (hartnn):

yup.

OpenStudy (xoterrixo13):

can u help with one m ore Find the midpoint of the segment with the given endpoints. D(1, 2) and E(–3, 6) -7,10 -2,2 -1,4 4,64

hartnn (hartnn):

take average of co-ordinates x= (x1+x2)/2 y=(y1+y2)/2

OpenStudy (xoterrixo13):

-1,4

hartnn (hartnn):

yes, correct.

OpenStudy (xoterrixo13):

Charlene made the sketch below in order to find the height x of a pole. She positioned a mirror on the ground so that she could see the reflection of the top of the pole. Her height, her distance from the mirror, and her line of sight to the mirror determine the smaller triangle. The pole’s height, its distance from the mirror, and the distance from the top of the pole to the mirror form a larger similar triangle. Find the height of the pole to the nearest tenth.

hartnn (hartnn):

new Q, new post please.